The gravitational force between two interacting particles spaced 2 cm apart is 6x10^-6 N. What should be the mass of one particle if other particle has a mass of 9 kg?


1
Expert's answer
2021-11-17T10:40:29-0500
"F=\dfrac{Gm_1m_2}{d^2},""m_1=\dfrac{Fd^2}{Gm_2}=\dfrac{6\times10^{-6}\ N\times(0.02\ m)^2}{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times9\ kg}=4\ kg."